OpenStudy (pokemon23):

One number is 5 more than another.The product is 14. Find the numbers.

OpenStudy (pokemon23):

\[x^2+5x=14\]

OpenStudy (chaise):

One number is five more than another. The product of the two numbers is 14. Find the two numbers? a+5=b a*b=14 5=b-a 14=a*b Two equations; two unknowns, correct me if I am wrong?

OpenStudy (anonymous):

is it 2

OpenStudy (anonymous):

and 7

Directrix (directrix):

Let x be one number Then, x + 5 is the other x(x+5) =14 x^2 + 5x = 14 x^2 + 5x – 14 = 0 (x + 7)(x-2) = 0 x = -7 or x = 2 Numbers: - 7 and -2 Or 2 and 7

OpenStudy (chaise):

a+5=b a*b=14 a(a+5)=14 a^2+5a=14 a^2+5a-14=0 (a-2)(a+7) a=2 a=-7 -7 is an imossible answer and therefore negated.

OpenStudy (chaise):

and then you sub back into the equation a+5=b 2+5=b b=7 the two numbers are a=2 and b=7

OpenStudy (pokemon23):

To solve a Quad equation 1)write equation in standfard form 2)Factor 3) Set each factor=0 Solve the resulting equation

Directrix (directrix):

@ Chaise Why - -7 is an imossible answer and therefore negated.

OpenStudy (chaise):

It's not really impossible - just negated because it yields the same answer. I guess if you really wanted to you could substitute a=-7, but for this case there is no real point.